# A Bidirectional Long Short Term Memory Approach for Infrastructure Health Monitoring Using On-board Vibration Response Authors: Reza Riahi Samani

### v1

@@ 5 Conclusions and future work
In this paper, we proposed a novel method for estimating and localizing multiple physical parameters of infrastructures using drive-by vibration responses, achieved through the development of an LSTM-BiLSTM network. Our methodology emphasizes the importance of considering both forward and backward temporal information through BiLSTM networks to enhance algorithms for monitoring infrastructure health conditions. The findings demonstrate that employing BiLSTM in the higher-level health condition estimation phase can significantly reduce MAPE and RMSE, even in the presence of noise. Moreover, the methodology highlights the potential of LSTM networks for extracting features from drive-by vibration signals. By accurately identifying the positions of vibration signals, our proposed framing approach improves the resolution of infrastructure monitoring to the beam level. Hence, the LSTM-BiLSTM model effectively estimates physical parameters at various components within the infrastructure beam levels, offering comprehensive insights into the health conditions of the systems and enhancing maintenance decision-making frameworks. In the case study analyzed, the results show that the method is effective in estimating the stiffness parameters of railpads and ballast, as well as in identifying reductions in track stiffness at the level of individual sleepers. The proposed BiLSTM networks significantly decrease the Mean Absolute Percentage Error (MAPE) and Root Mean Square Error (RMSE) by nearly 50%, achieving a MAPE of 1.7% for railpad stiffness and 0.7% for ballast stiffness, even in noisy environments.
Future research will focus on assessing the proposed model’s performance with other benchmarks in the literature to provide more robust conclusions on the proposed BiLSTM networks. Moreover, the feature extraction phase requires further exploration of advanced CNN and LSTM networks, alongside other preprocessing techniques for vibration response signals. The proposed framing approach requires high-accuracy signal localization, which can be challenging under certain operational conditions. In this regard, leveraging domain knowledge concerning the number and location of beam nodes in the infrastructure holds promise for addressing the location approximation challenge.

### v2

@@ 5 Conclusions and future work
In this paper, we proposed a novel method for estimating and localizing multiple physical parameters of infrastructures using drive-by vibration responses, achieved through the development of an LSTM-BiLSTM network. Our methodology emphasizes the importance of considering both forward and backward temporal information through BiLSTM networks to enhance algorithms for monitoring infrastructure health conditions. The findings demonstrate that employing BiLSTM in the higher-level health condition estimation phase can significantly reduce MAPE and RMSE, even in the presence of noise. Moreover, the methodology highlights the potential of LSTM networks for extracting features from drive-by vibration signals. By accurately identifying the positions of vibration signals, our proposed framing approach improves the resolution of infrastructure monitoring to the beam level. Hence, the LSTM-BiLSTM model effectively estimates physical parameters at various components within the infrastructure beam levels, offering comprehensive insights into the health conditions of the systems and enhancing maintenance decision-making frameworks. In the case study analyzed, the results show that the method is effective in estimating the stiffness parameters of railpads and ballast, as well as in identifying track stiffness reductions at individual sleepers. The proposed BiLSTM networks significantly decrease the Mean Absolute Percentage Error (MAPE) and Root Mean Square Error (RMSE) by nearly 50%, achieving a MAPE of 1.7% for railpad stiffness and 0.7% for ballast stiffness, even in noisy environments.
Future research will focus on three main research directions. First, the proposed hybrid model architecture requires further benchmarking against state-of-the-art methodologies. This includes considering advanced CNN and LSTM networks alongside other machine-learning methods. Second, future research should validate the performance of the proposed models based on field measurements and under real-world conditions. This further requires developing deep learning models that are robust to real-world noisy environments. Thirdly, the proposed framing approach requires high-accuracy signal localization, which can be challenging under certain operational conditions. In this regard, leveraging domain knowledge concerning the number of beam nodes in certain infrastructure lengths holds promise for addressing the signal localization challenge.